# Reconstructed Electronic Structure in 2D vdW 1T-Ta for Quantum Sensing and Information Science

### v2

@@ Abstract
American Chemical Society, LaTeXIntroduction- Transition metal dichalcogenides (TMDs), a class of two-dimensional (2D) layered materials, possess have a range of electronic properties from insulating to semiconducting, metallic and even superconducting 1, 2, 3. It is pretty interesting to understand the driving mechanism behind the formation of the charge density waves (CDWs) in these systems, which is still unclear to address 4, 5. In general understanding, it is expected that perfection in Fermi-surface nesting (FSN) cause divergence in the electronic susceptibility. This process may cause the existence of CDW phases in this system. Meanwhile, scattering behaviour affect the atomic coherence and diminish this coherency when this material goes to a correlated electron state and vice versa. Thus, it is absolutely needed to persist such correlated electronic state in system to achieve long term CDW phase stability, where large and sharp interfaces in materials support this feature 6, 7. Considering these facts, 2D materials possess atomically large and sharp interfaces helping to hold robustly the correlated states over long distances with reduced scattering 8. Moreover, the graphene family and related 2D heterostructures support the observation of strong coupling effects (i.e. superconductivity, magnetism and spin-orbit coupling) at low dimensional length scale 9, 10.
@@
Here, 1T-Ta is a TMD compound that exhibits a rich variety of CDW phases. The CDW in 1T-Ta is known to undergo several structural transitions, leading to the formation of commensurate CDW (CCDW) states. The understanding of the electronic structure and photoemission matrix element effects in these CDW phases is crucial for unraveling their complex behavior. However, there is increasing high questions raised to understand purely the FSN effect and electronic origins of the instability due to the CDW phase in layered quasi 2D CDW materials. In this line, first-principles reports on layered CDW materials (such as 1T-Ta, 1T-Ta, and 2H-Nb) show correlation-less between CDW ordering vector and strong nesting vectors that supports the role of electron-phonon interactions in the CDW formation. Meanwhile, the CDW phase can be extensively investigated as a another many-body quantum state due to electron-electron correlations 11, like on-site Coulomb repulsion and anti-ferromagnetic short-range spin-exchange interactions, predicted in various correlated insulators and unconventional superconductors 12. Previously, there is an instance of CDW phases observed in 2D TMDs 13, 14. It is challenging to study the chiral effect between normal and CDW phase of the same material including the role of scattering surfaces. Present work is considered for the 1T-Ta as a prototypical TMD having rich CDWs phase diagram 15. Using first-principles calculations, FSN effect, van Hove singularities and electron-phonon coupling are examined on the structural and electronic properties of the system 1, 16, 17. We further simulate the photoemission electronic spectra to understand the matrix element effects on the system.
@@
Reconstructed Electronic Structure in 1T-Ta: The electronic structure of 1T-Ta is significantly modified in the presence of CDW phases. Several experimental techniques, such as angle-resolved photoemission spectroscopy (ARPES), scanning tunneling microscopy/spectroscopy (STM/STS), and X-ray scattering, have been employed to probe the electronic structure. ARPES measurements revealed the formation of an energy gap in the CDW state, indicating the opening of a partial band gap due to the reconstruction of the electronic structure 18. The band dispersion and the Fermi surface topology were found to be strongly modified in the CDW phase. STM/STS studies provided further insights into the CDW-induced electronic reconstruction and the emergence of localized states within the gap 19.
@@
Methodology Set-ups- The full cell geometry optimization and electronic structure calculations of the 1T-Ta system are performed using density functional theory (DFT) simulations. Here, projector augmented wave (PAW) 20, 21 formalism is considered during calculations with the Quantum ESPRESSO Package 22, 23, 24, 25. The Perdew-Burke-Ernzerhof (PBE) generalized-gradient approximation (GGA), specially for solids (PBESol) 26 is generally used to describe the exchange and correlation functional. The CDW structures are constructed in a supercell of the 1T-Ta starting from the experimental periodic lattice distortion (PLD) positions of atoms taken from Ref. 27. A 20 vacuum gap is included out-of-plane along c axis to avoid periodic image interaction. Here, full cell volume relaxation is performed with Methfessel-Paxton smearing () of 0.2 eV for all sets of calculations. All the geometric structures are fully relaxed until the force and energy tolerance convergence on each atom become 0.002 eV/ and eV, respectively. A set of K-point mesh is chosen to appropriately sample the Brillouin zone with respect to the primitive cell for the computation of the structure relaxations and the self consistent electronic densities and a denser, mesh for the computation of the densities of states. We set 580 eV as the cutoff energy and rev-vdW-DF2 28, 29 vdW correction for the calculations. The atoms in the CDW phase are fully relaxed while the lattice parameters are fixed at the values of the pristine phase. The tetrahedron method is employed in calculating the density of states and total energy. The phonon dispersion is calculated by using the finite displacement method using the PHONOPY code 30. Hamiltonian of tight-binding basis is used for FSN, photo-emmision matrix element properties based on the maximally localized Wannier functions (MLWFs) of WANNIER90 package 31. Moreover, simulate angle-resolved photoemission matrix elements are compared together to check the compatibility 32.
@@
Tight-binding model calculations- We follow tight-binding model to present the reconstructed electronic structures of the 1T-Ta system with commensurate CDW locked in star-of-David cluster form. This model uses qualitatively the primary results of band structures in normal and CCDW phases. The Hamiltonian is formed by the Slater-Koster (SK) method 33. We considered , , and orbitals of Ta atoms as the basis due to their dominant band structure near Fermi level () during the calculations 34. It is noticed that in-plane displacements of Ta atoms are generally twice larger than the out-of-plane displacements of S atoms, while the system enters CCDW phase 35. In case of high temperature normal phase calculations, one Ta atom is present with two S atoms in the unit cell having a 6*6 square matrix, whereas 13 Ta atoms and 26 S atoms are present in the superlattice unit cell of the CCDW phase having matrix (in case of spin orbit coupling). We can obtain the energy eigen values along with the band dispersions and constant-energy contours by diagonalizing the Hamiltonian matrix. We obtain the tight-binding model parameters from literature 34 and use in this calculations. We scale the SK hopping integrals with respect to the inverse cubic Ta-Ta distance to fit into the simulated photo-emission matrix elements.
@@
Results- Bulk 1T-Ta takes the commensurate CDW phase below 170K. Thus, it is difficult to obtain the lattice parameters of bulk CCDW 1T-Ta phase experimentally 36. We start with the full cell relaxation of the atomic positions and respective lattice parameters in the superlattice unit cell taking 26 S atoms and 13 Ta atoms in bulk configuration. The optimized superlattice parameter is 11.98 and Ta-Ta shortest bonds inside the SOD pattern is 3.135 (compared with 3.317 in the high symmetry phase).
@@
Commensurate Charge Density Wave States: The CDW in 1T-Ta can exist in various commensurate phases, characterized by different periodicity. These phases include 4a, 4a+b, 5a, 6a, and more complex superlattice structures. Each commensurate CDW phase exhibits distinct electronic properties and ordering patterns. In this regard, x-ray scattering experiments revealed the formation of a 5a CDW phase with a periodicity of five times the lattice constant, leading to the formation of an unusual quasi-periodic superlattice 37. The CDW-induced structural reconstruction helps to observe the electronic modulations. Our obtained band structures for bulk and monolayer CCDW 1T-Ta phases are shown in Figure 2 and Figure 3, respectively. The resulting minimizing forces and stress of structure relaxation are reconstructed as SOD phase. It shows the unit cell of the periodic lattice distortions (PLD) and SOD clusters in the Ta plane (shown in Figure 4). Here, the commensurate phase of the CDW phase can only be computationally accessible for periodic DFT simulations, thus we focus of the commensurate-CDW (CCDW) phase. Moreover, it is noticed that the superlattice unit cell contains maximum three types of non-equivalent Ta atoms. Here, a=+3 and b=4- and c= are the three lattice vectors in the CCDW phase that form the superlattice with these in-plane basis vectors. Here, , and parameters follow the lattice vectors as per the hexagonal 1T-Cd type structure.
@@
Fermi surface nesting (FSN): The photoemission matrix element effects can influence the observed Fermi surface nesting (FSN) in 1T-Ta with commensurate CDW phases 37. Here, FSN refers to the alignment of Fermi surface portions that can facilitate the formation of CDWs through the scattering of electrons between these portions. The photoemission matrix element effects modify the photoemission spectra and, consequently, affect the observed the FSN (refer to Figure 3). In 1T-Ta, the CDWs introduce an intricate reconstruction of the electronic structure, leading to the opening of partial band gaps and modifications in the Fermi surface topology (i.e. the Fermi surface states, FSS). The photoemission matrix elements, which depend on the electronic and lattice structure, can enhance or suppress specific spectral features in the photoemission spectra 38, 39, relevant near the CDW-induced band gaps contributing the observed nesting behavior.
@@
Periodic lattice distortions in CCDW phase- The normal phase 1T-Ta system takes high symmetry state P-3m1 (Space group 164). We performed full wanniersization of Ta d-orbitals using WANNIER90 and check its convergence with DFT band structure for high symmetry phase. Fermi surface has been plotted from the WANNIER90 results and visualized using XCrysDen software 40 (shown in Figure S1). In the superlattice unit cell, total 13 Tantalum (Ta) atoms and 26 Sulfur (S) atoms present to form the structure. To obtain the commensurate CDW phase from this superlattice undistorted unit cell, we induce a periodic lattice distortions into all the Ta atoms centering one Ta atoms. A random displacement (within 1-3) to all 12 Ta atoms has been performed in-plane coordinates following a full cell relaxation. After relaxation, we get the distorted CCDW phase of 1T-Ta superlattice unit cell (shown in Figure S2). We checked both bulk ((shown in Figure S3)) and monolayer ((shown in Figure S4)) cases of CCDW phase and notice the up-gradation in electronic band structures and formation of the Star of David (SOD) clusters. The underneath mechanism of FSN in 1T-Ta with commensurate CDW phases is intimately linked to the PLD that accompany the CDW formation. These distortions are crucial for the emergence of nesting conditions, enabling the scattering of electrons between Fermi surface portions and facilitating the formation of CDWs. The specific details of the nesting mechanism, including the nesting vectors and the resulting CDW wave vectors, depend on the characteristics of the lattice distortions and the resulting band folding in the material. Understanding and controlling these mechanisms are crucial for exploring the rich physics and potential applications of CDW materials like 1T-Ta.
@@
Discussions- One possible explanation for the primary goal of this chirality may be given by the superlattice reconstruction. One symmetry k-point () is present in the first Brillouin zone (BZ) represented by an effective wave vector in case of normal undistorted unit cell. Similarly, the CCDW distorted phase takes additional two cases of symmetry k-points, showing perfect reduction in the BZ wave vectors. This proper reduction in BZ wave vectors are framed as dashed smaller hexagons of black coloured for reduced BZ of distorted CCDW phase and solid big hexagon of red coloured as BZ of undistorted normal phase, having 13. rotation between normal and CCDW phase BZ (shown in the inset of Figure 1). Moreover, elliptical Fermi surface pockets are marked in normal phase to understand the nesting and rotational directions of the reconstructed CCDW phase. Our first principle simulation result is consistent with photoemission matrix elements data, where one Ta-5d band crosses the Fermi level (), forming 06 oval-type packets surrounding the BZ boundaries. Meanwhile, the CDW potential retains the BZ in the CCDW phase and the Fermi surface reconstructs into multiple reduced spots with (×)R13. periodicity. Thus, the three types of BZ points may be responsible for the possible polarization directions 41, 42, 43.
@@
We can notice that Ta atoms are distorted towards the central Ta atoms of the SOD cluster (Figure 2). We find the a-b and a-c interatomic distances that compressed by 4.7 (4.5 expt. value) and 3.2 (2.9 expt. value), respectively which is in close agreement with experimental report 27. Mostly, the Ta atoms distort in-plane direction, so S atoms displace towards out-plane direction (Figure 1, bottom right panel) due to the Ta interatomic compression. Figure LABEL:figure-3 shows the band structure of CCDW phase that is different from the bulk phase. Here, an in-plane band gap opens at the Fermi level, which is the most noticeable fact. Here, Ta d-bands form strongly localized bands up to -1.0 eV the valence band below Fermi level. It is observed that the uppermost p-bands are localized near to -1.48 eV, which is in agreement with experiment 44, 45. This supports the projected density of states and electronically driven instability via Fermi surface nesting in the bulk CCDW phase.

### v3

@@ Abstract
American Chemical Society, LaTeXIntroduction- Transition metal dichalcogenides (TMDs), a class of two-dimensional (2D) layered materials, possess have a range of electronic properties from insulating to semiconducting, metallic and even superconducting PhysRevLett.117.106801, PhysRevLett.125.186401, PhysRevLett.125.165302. It is pretty interesting to understand the driving mechanism behind the formation of the charge density waves (CDWs) in these systems, which is still unclear to address PhysRevLett.100.036804, PhysRevLett.129.156401. In general understanding, it is expected that perfection in Fermi-surface nesting (FSN) cause divergence in the electronic susceptibility. This process may cause the existence of CDW phases in this system. Meanwhile, scattering behaviour affect the atomic coherence and diminish this coherency when this material goes to a correlated electron state and vice versa. Thus, it is absolutely needed to persist such correlated electronic state in system to achieve long term CDW phase stability, where large and sharp interfaces in materials support this feature PhysRevLett.130.226401, PhysRevLett.120.037002. Considering these facts, 2D materials possess atomically large and sharp interfaces helping to hold robustly the correlated states over long distances with reduced scattering PhysRevLett.122.086402. Moreover, the graphene family and related 2D heterostructures support the observation of strong coupling effects (i.e. superconductivity, magnetism and spin-orbit coupling) at low dimensional length scale PhysRevB.77.033306, Behera_2021.
@@
Here, 1T-Ta is a TMD compound that exhibits a rich variety of CDW phases. The CDW in 1T-Ta is known to undergo several structural transitions, leading to the formation of commensurate CDW (CCDW) states. The understanding of the electronic structure and photoemission matrix element effects in these CDW phases is crucial for unraveling their complex behavior. However, there is increasing high questions raised to understand purely the FSN effect and electronic origins of the instability due to the CDW phase in layered quasi 2D CDW materials. In this line, first-principles reports on layered CDW materials (such as 1T-Ta, 1T-Ta, and 2H-Nb) show correlation-less between CDW ordering vector and strong nesting vectors that supports the role of electron-phonon interactions in the CDW formation. Meanwhile, the CDW phase can be extensively investigated as a another many-body quantum state due to electron-electron correlations PhysRevLett.106.136809, like on-site Coulomb repulsion and anti-ferromagnetic short-range spin-exchange interactions, predicted in various correlated insulators and unconventional superconductors Keimer2015. Previously, there is an instance of CDW phases observed in 2D TMDs PhysRevB.107.L201109, C8CP04684K. It is challenging to study the chiral effect between normal and CDW phase of the same material including the role of scattering surfaces. Present work is considered for the 1T-Ta as a prototypical TMD having rich CDWs phase diagram PhysRevX.12.021064. Using first-principles calculations, FSN effect, van Hove singularities and electron-phonon coupling are examined on the structural and electronic properties of the system PhysRevLett.117.106801, PhysRevLett.121.027003, PhysRevLett.99.216404. We further simulate the photoemission electronic spectra to understand the matrix element effects on the system.
@@
Reconstructed Electronic Structure in 1T-Ta: The electronic structure of 1T-Ta is significantly modified in the presence of CDW phases. Several experimental techniques, such as angle-resolved photoemission spectroscopy (ARPES), scanning tunneling microscopy/spectroscopy (STM/STS), and X-ray scattering, have been employed to probe the electronic structure. ARPES measurements revealed the formation of an energy gap in the CDW state, indicating the opening of a partial band gap due to the reconstruction of the electronic structure Chen2015. The band dispersion and the Fermi surface topology were found to be strongly modified in the CDW phase. STM/STS studies provided further insights into the CDW-induced electronic reconstruction and the emergence of localized states within the gap pnas.1211387110.
@@
Methodology Set-ups- The full cell geometry optimization and electronic structure calculations of the 1T-Ta system are performed using density functional theory (DFT) simulations. Here, projector augmented wave (PAW) PhysRevB.50.17953, PhysRevB.59.1758 formalism is considered during calculations with the Quantum ESPRESSO Package PhysRevB.47.558, PhysRevB.49.14251, PhysRevB.54.11169, KRESSE199615. The Perdew-Burke-Ernzerhof (PBE) generalized-gradient approximation (GGA), specially for solids (PBESol) PhysRevB.79.155107 is generally used to describe the exchange and correlation functional. The CDW structures are constructed in a supercell of the 1T-Ta starting from the experimental periodic lattice distortion (PLD) positions of atoms taken from Ref. PhysRevB.56.13757. A 20 vacuum gap is included out-of-plane along c axis to avoid periodic image interaction. Here, full cell volume relaxation is performed with Methfessel-Paxton smearing () of 0.2 eV for all sets of calculations. All the geometric structures are fully relaxed until the force and energy tolerance convergence on each atom become 0.002 eV/ and eV, respectively. A set of K-point mesh is chosen to appropriately sample the Brillouin zone with respect to the primitive cell for the computation of the structure relaxations and the self consistent electronic densities and a denser, mesh for the computation of the densities of states. We set 580 eV as the cutoff energy and rev-vdW-DF2 dft-d3, dft-d3-1 vdW correction for the calculations. The atoms in the CDW phase are fully relaxed while the lattice parameters are fixed at the values of the pristine phase. The tetrahedron method is employed in calculating the density of states and total energy. The phonon dispersion is calculated by using the finite displacement method using the PHONOPY code phonopy. Hamiltonian of tight-binding basis is used for FSN, photo-emmision matrix element properties based on the maximally localized Wannier functions (MLWFs) of WANNIER90 package MOSTOFI2008685. Moreover, simulate angle-resolved photoemission matrix elements are compared together to check the compatibility PhysRevLett.122.127001.
@@
Tight-binding model calculations- We follow tight-binding model to present the reconstructed electronic structures of the 1T-Ta system with commensurate CDW locked in star-of-David cluster form. This model uses qualitatively the primary results of band structures in normal and CCDW phases. The Hamiltonian is formed by the Slater-Koster (SK) method SUGAI1981405. We considered , , and orbitals of Ta atoms as the basis due to their dominant band structure near Fermi level () during the calculations PhysRevB.73.073106. It is noticed that in-plane displacements of Ta atoms are generally twice larger than the out-of-plane displacements of S atoms, while the system enters CCDW phase Rossnagel2011. In case of high temperature normal phase calculations, one Ta atom is present with two S atoms in the unit cell having a 6*6 square matrix, whereas 13 Ta atoms and 26 S atoms are present in the superlattice unit cell of the CCDW phase having matrix (in case of spin orbit coupling). We can obtain the energy eigen values along with the band dispersions and constant-energy contours by diagonalizing the Hamiltonian matrix. We obtain the tight-binding model parameters from literature PhysRevB.73.073106 and use in this calculations. We scale the SK hopping integrals with respect to the inverse cubic Ta-Ta distance to fit into the simulated photo-emission matrix elements.
@@
Results- Bulk 1T-Ta takes the commensurate CDW phase below 170K. Thus, it is difficult to obtain the lattice parameters of bulk CCDW 1T-Ta phase experimentally PhysRevB.105.L081106. We start with the full cell relaxation of the atomic positions and respective lattice parameters in the superlattice unit cell taking 26 S atoms and 13 Ta atoms in bulk configuration. The optimized superlattice parameter is 11.98 and Ta-Ta shortest bonds inside the SOD pattern is 3.135 (compared with 3.317 in the high symmetry phase).
@@
Commensurate Charge Density Wave States: The CDW in 1T-Ta can exist in various commensurate phases, characterized by different periodicity. These phases include 4a, 4a+b, 5a, 6a, and more complex superlattice structures. Each commensurate CDW phase exhibits distinct electronic properties and ordering patterns. In this regard, x-ray scattering experiments revealed the formation of a 5a CDW phase with a periodicity of five times the lattice constant, leading to the formation of an unusual quasi-periodic superlattice PhysRevB.107.165115. The CDW-induced structural reconstruction helps to observe the electronic modulations. Our obtained band structures for bulk and monolayer CCDW 1T-Ta phases are shown in Figure 2 and Figure 3, respectively. The resulting minimizing forces and stress of structure relaxation are reconstructed as SOD phase. It shows the unit cell of the periodic lattice distortions (PLD) and SOD clusters in the Ta plane (shown in Figure 4). Here, the commensurate phase of the CDW phase can only be computationally accessible for periodic DFT simulations, thus we focus of the commensurate-CDW (CCDW) phase. Moreover, it is noticed that the superlattice unit cell contains maximum three types of non-equivalent Ta atoms. Here, a=+3 and b=4- and c= are the three lattice vectors in the CCDW phase that form the superlattice with these in-plane basis vectors. Here, , and parameters follow the lattice vectors as per the hexagonal 1T-Cd type structure.
@@
Fermi surface nesting (FSN): The photoemission matrix element effects can influence the observed Fermi surface nesting (FSN) in 1T-Ta with commensurate CDW phases PhysRevB.107.165115. Here, FSN refers to the alignment of Fermi surface portions that can facilitate the formation of CDWs through the scattering of electrons between these portions. The photoemission matrix element effects modify the photoemission spectra and, consequently, affect the observed the FSN (refer to Figure 3). In 1T-Ta, the CDWs introduce an intricate reconstruction of the electronic structure, leading to the opening of partial band gaps and modifications in the Fermi surface topology (i.e. the Fermi surface states, FSS). The photoemission matrix elements, which depend on the electronic and lattice structure, can enhance or suppress specific spectral features in the photoemission spectra PhysRevB.77.235104, PhysRevB.105.235104, relevant near the CDW-induced band gaps contributing the observed nesting behavior.
@@
Periodic lattice distortions in CCDW phase- The normal phase 1T-Ta system takes high symmetry state P-3m1 (Space group 164). We performed full wanniersization of Ta d-orbitals using WANNIER90 and check its convergence with DFT band structure for high symmetry phase. Fermi surface has been plotted from the WANNIER90 results and visualized using XCrysDen software KOKALJ1999176 (shown in Figure S1). In the superlattice unit cell, total 13 Tantalum (Ta) atoms and 26 Sulfur (S) atoms present to form the structure. To obtain the commensurate CDW phase from this superlattice undistorted unit cell, we induce a periodic lattice distortions into all the Ta atoms centering one Ta atoms. A random displacement (within 1-3) to all 12 Ta atoms has been performed in-plane coordinates following a full cell relaxation. After relaxation, we get the distorted CCDW phase of 1T-Ta superlattice unit cell (shown in Figure S2). We checked both bulk ((shown in Figure S3)) and monolayer ((shown in Figure S4)) cases of CCDW phase and notice the up-gradation in electronic band structures and formation of the Star of David (SOD) clusters. The underneath mechanism of FSN in 1T-Ta with commensurate CDW phases is intimately linked to the PLD that accompany the CDW formation. These distortions are crucial for the emergence of nesting conditions, enabling the scattering of electrons between Fermi surface portions and facilitating the formation of CDWs. The specific details of the nesting mechanism, including the nesting vectors and the resulting CDW wave vectors, depend on the characteristics of the lattice distortions and the resulting band folding in the material. Understanding and controlling these mechanisms are crucial for exploring the rich physics and potential applications of CDW materials like 1T-Ta.
@@
Discussions- One possible explanation for the primary goal of this chirality may be given by the superlattice reconstruction. One symmetry k-point () is present in the first Brillouin zone (BZ) represented by an effective wave vector in case of normal undistorted unit cell. Similarly, the CCDW distorted phase takes additional two cases of symmetry k-points, showing perfect reduction in the BZ wave vectors. This proper reduction in BZ wave vectors are framed as dashed smaller hexagons of black coloured for reduced BZ of distorted CCDW phase and solid big hexagon of red coloured as BZ of undistorted normal phase, having 13. rotation between normal and CCDW phase BZ (shown in the inset of Figure 1). Moreover, elliptical Fermi surface pockets are marked in normal phase to understand the nesting and rotational directions of the reconstructed CCDW phase. Our first principle simulation result is consistent with photoemission matrix elements data, where one Ta-5d band crosses the Fermi level (), forming 06 oval-type packets surrounding the BZ boundaries. Meanwhile, the CDW potential retains the BZ in the CCDW phase and the Fermi surface reconstructs into multiple reduced spots with (×)R13. periodicity. Thus, the three types of BZ points may be responsible for the possible polarization directions Wang2020, acs.nanolett.1c04990, PhysRevB.96.125138.
@@
We can notice that Ta atoms are distorted towards the central Ta atoms of the SOD cluster (Figure 2). We find the a-b and a-c interatomic distances that compressed by 4.7 (4.5 expt. value) and 3.2 (2.9 expt. value), respectively which is in close agreement with experimental report PhysRevB.56.13757. Mostly, the Ta atoms distort in-plane direction, so S atoms displace towards out-plane direction (Figure 1, bottom right panel) due to the Ta interatomic compression. Figure LABEL:figure-3 shows the band structure of CCDW phase that is different from the bulk phase. Here, an in-plane band gap opens at the Fermi level, which is the most noticeable fact. Here, Ta d-bands form strongly localized bands up to -1.0 eV the valence band below Fermi level. It is observed that the uppermost p-bands are localized near to -1.48 eV, which is in agreement with experiment Manzke_1989, PhysRevB.89.155137. This supports the projected density of states and electronically driven instability via Fermi surface nesting in the bulk CCDW phase.